5 whole 1/6 ÷ 9/2 is equal to: (a) 31/6 (b) 1/27 (c) (d) 31/27

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Option (d) is correct here.
First we have to convert the mixed fraction into improper fraction 5 whole 1/6 = 31/6 Then
= 31/6 ÷ 9/2
= (31/6)/ (9/2)
= (31/6) × (2/9)
= (31/3) × (1/9)
= 31/27


Thanks